OpenStudy (anonymous):

How do you solve for x^{4}+x^{5}-1=0

OpenStudy (anonymous):

You don't solve these kind of equations, they're too hard.

OpenStudy (anonymous):

You can approximate x though.

OpenStudy (anonymous):

graphically or?

OpenStudy (anonymous):

That's one way to do it, or you ask a computer.

OpenStudy (anonymous):

Bah too hard! :D

OpenStudy (anonymous):

Quadradic: \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] for in the ax\(^2\) + bx + c = 0 form

OpenStudy (anonymous):

strip the coefficients off, and don't forget that if it's like this you need to pull the negatives too: x^2-x-1=0 x^2+(-1)x+(-1)=0 a = 1 b = -1 c = -1

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Yep you have five answers and only one of them is a real root

OpenStudy (anonymous):

The equation has one real root and 4 imaginary roots.

OpenStudy (anonymous):

Make a substitution if you want to use the quadratic, at first I thought that was a x^2

OpenStudy (anonymous):

So it's messy, yes. I'd agree with you there @Thomas9

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

You cannot use the quadratic formula here.

OpenStudy (anonymous):

exactly quadratic cannot work...

OpenStudy (anonymous):

Two words: Fractional exponent.

OpenStudy (anonymous):

Agree?

OpenStudy (anonymous):

somehow

OpenStudy (anonymous):

Eliassaab i think you r so so right there...quadratic now am convinced it cant..

OpenStudy (anonymous):

This question has been answered but not I'm going to look into this more, maybe you can't do a fractional substitution with a "quintic" polynomial... Hmm... This is along the lines of what I was thinking \[ax^4+bx^2+c\] let u = x\(^2\) \[au^2+bu+c\] \[ax^5+bx^{2.5}+c\] let u = x\(^{2.5}\)=\(\sqrt{x^5}\) \[au^2+bu+c\]

OpenStudy (anonymous):